OpenStudy (anonymous):

how do u write 6000 in scientific notation please helpp!

OpenStudy (math&ing001):

\[6000 = 6.10^{3}\]

OpenStudy (igreen):

Actually it's: \(6 \times 10^3\). The exponent 3, is the number of zeroes.

OpenStudy (math&ing001):

That's exactly what I said @iGreen !

OpenStudy (anonymous):

how about 5003

OpenStudy (igreen):

You wrote \(6.10^3\) which looks like the decimal 6.10 with an exponent of 3.

OpenStudy (math&ing001):

Nope . means multiplication

OpenStudy (igreen):

That would be \(5.003 \times 10^3\)..make sure you include the 3 in your decimal.

OpenStudy (igreen):

Yes, but you should use a multiplication symbol like \(\times\) or \(·\), and not a period.

OpenStudy (math&ing001):

In physics we use . instead of x

OpenStudy (igreen):

Then use the centered dot, not a period.

OpenStudy (igreen):

(⋅) not (.)

OpenStudy (anonymous):

Guys.. guys.. guys.. No need to argue.

OpenStudy (math&ing001):

It's pathetic to argue over whether a dot is centered or not, you got the meaning and so did he. That's what's important.

OpenStudy (igreen):

I didn't get the meaning till you told me, and he never said he got the meaning.

OpenStudy (igreen):

It looks like a decimal..